Orange, Inc. has identified the following cost drivers for its expected overhead costs for the year:

Cost Pools Budgeted Cost Cost Driver Cost Driver Level
Setup $ 220,000 Number of setups 1,100
Ordering 110,000 Number of orders 5,500
Maintenance 275,000 Machine hours 27,500
Power 55,000 Kilowatt hours 55,000

Total direct labor hours budgeted = 11,000 hours.

The following data applies to Product X, one of the products completed during the year.

Direct materials $ 5,500
Direct labor $ 6,600
Units completed 550
Direct labor hours 220
Number of setups 22
Number of orders 44
Machine hours 275
Kilowatt hours 550

If the activity-based cost drivers are used to allocate overhead cost, the total overhead cost of Product X will be:

Multiple Choice

  • $8,520.

  • $8,580.

  • $11,000.

  • $13,200.

  • $9,580.
Answer #1

Answer

--Correct Answer = Option #2: $ 8580

Cost pool Overhead rate Activity driver for Product X Overhead cost allocated
Setup 200 per setup 22 setup $4,400
Ordering 20 per orders 44 orders $880
Maintenance 10 per MHs 275 MHs $2,750
Power 1 per KHs 550 KHs $550
$8,580 Answer

--Calculation

Cost pool Overhead rate Activity driver for Product X Overhead cost allocated
Setup =220000/1100 per setup 22 setup =200*22
Ordering =110000/5500 per orders 44 orders =20*44
Maintenance =275000/27500 per MHs 275 MHs =10*275
Power =55000/55000 per KHs 550 KHs =1*550